Find the general solution of the following differential equations using method of undetermined coefficients


(i) y''−7y'+10y =20e^5x −10

(ii) y''−2y'= x^2 +5x−2

(iii) y''+ 9y'+14y = 20+e^2x +e^−2x

(iv) y''−2y'+y =e^x+x^2

(v) y''−y'=cosx+5.


1
Expert's answer
2022-03-03T07:24:30-0500

(i)

Corresponding homogeneous differential equation


"y''\u22127y'+10y =0"

Characteristic (auxiliary) equation


"r^2-7r+10=0"

"r_1=2,r_2=5"

The general solution of the homogeneous differential equation is


"y_h=c_1e^{2x}+c_2e^{5x}"

Find the particular solution of the non homogeneous differential equation


"y_p=Axe^{5x}+B"

"y_p'=5Axe^{5x}+Ae^{5x}"

"y_p''=25Axe^{5x}+10Ae^{5x}"

Substitute


"25Axe^{5x}+10Ae^{5x}-35Axe^{5x}-7Ae^{5x}"

"+10Axe^{5x}+10B=20e^{5x}-10"

"A=20\/3, B=-1"

"y_p=\dfrac{20}{3}xe^{5x}-1"

The general solution of the non homogeneous differential equation is


"y=c_1e^{2x}+c_2e^{5x}+\dfrac{20}{3}xe^{5x}-1"



(ii)

Corresponding homogeneous differential equation


"y''\u22122y' =0"

Characteristic (auxiliary) equation


"r^2-2r=0"

"r_1=0,r_2=2"

The general solution of the homogeneous differential equation is


"y_h=c_1+c_2e^{2x}"

Find the particular solution of the non homogeneous differential equation


"y_p=Ax^3+Bx^2+Cx"

"y_p'=3Ax^2+2Bx+C"

"y_p''=6Ax+2B"

Substitute


"6Ax+2B-6Ax^2-4Bx-2C"

"=x^2+5x-2"

"-6A=1"

"6A-4B=5"

"2B-2C=-2"

"A=-1\/6, B=-3\/2, C=-1\/2"

"y_p=-\dfrac{1}{6}x^3-\dfrac{3}{2}x^2-\dfrac{1}{2}x"

The general solution of the non homogeneous differential equation is


"y=c_1+c_2e^{2x}-\dfrac{1}{6}x^3-\dfrac{3}{2}x^2-\dfrac{1}{2}x"



(iii)

Corresponding homogeneous differential equation


"y''+9y'+14y =0"

Characteristic (auxiliary) equation


"r^2+9r+14=0"

"r_1=-7,r_2=-2"

The general solution of the homogeneous differential equation is


"y_h=c_1e^{-7x}+c_2e^{-2x}"

Find the particular solution of the non homogeneous differential equation


"y_p=Ae^{2x}+Bxe^{-2x}+C"




"y_p'=2Ae^{2x}-2Bxe^{-2x}+Be^{-2x}"

"y_p''=4Ae^{2x}+4Bxe^{-2x}-4Be^{-2x}"

Substitute


"4Ae^{2x}+4Bxe^{-2x}-4Be^{-2x}"

"+18Ae^{2x}-18Bxe^{-2x}+9Be^{-2x}"

"+14Ae^{2x}+14Bxe^{-2x}+14C"

"=20+e^{2x} +e^{\u22122x}"

"36A=1"

"5B=1"

"14C=20"

"y_p=\dfrac{1}{36}e^{2x}+\dfrac{1}{5}xe^{-2x}+\dfrac{10}{7}"

The general solution of the non homogeneous differential equation is


"y=c_1e^{-7x}+c_2e^{-2x}+\dfrac{1}{36}e^{2x}+\dfrac{1}{5}xe^{-2x}+\dfrac{10}{7}"



(iv)

Corresponding homogeneous differential equation


"y''\u22122y'+y =0"

Characteristic (auxiliary) equation


"r^2-2r+1=0"

"r_1=r_2=1"

The general solution of the homogeneous differential equation is


"y_h=c_1e^x+c_2xe^x"

Find the particular solution of the non homogeneous differential equation


"y_p=Ax^2e^{x}+Bx^2+Cx+D"




"y_p'=Ax^2e^x+2Axe^x+2Bx+C"

"y_p''=Ax^2e^x+4Axe^x+2Ae^x+2B"

Substitute


"Ax^2e^x+4Axe^x+2Ae^x+2B"

"-2Ax^2e^x-4Axe^x-4Bx-2C"

"+Ax^2e^{x}+Bx^2+Cx+D"

"=e^x+x^2"

"2A=1"

"B=1"

"C=4"

"D=6"

"y_p=\dfrac{1}{2}x^2e^{x}-x^2+4x+6"

The general solution of the non homogeneous differential equation is


"y=c_1e^x+c_2xe^x+\dfrac{1}{2}x^2e^{x}-x^2+4x+6"



(v)

Corresponding homogeneous differential equation


"y''\u2212y' =0"

Characteristic (auxiliary) equation


"r^2-r=0"

"r_1=0,r_2=1"

The general solution of the homogeneous differential equation is


"y_h=c_1+c_2e^{x}"

Find the particular solution of the non homogeneous differential equation


"y_p=A\cos x+B\sin x+Cx"

y''−y'=cosx+5.

"y_p'=-A\sin x+B\cos x+C"

"y_p''=-A\cos x-B\sin x"

Substitute

"-A\cos x-B\sin x+A\sin x+B\cos x-C"

"=\cos x+5"

"A=-1\/2, B=-1\/2, C=-5"

"y_p=-\dfrac{1}{2}\cos x-\dfrac{1}{2}\sin x-5x"

The general solution of the non homogeneous differential equation is


"y=c_1+c_2e^{x}-\dfrac{1}{2}\cos x-\dfrac{1}{2}\sin x-5x"